Calculate the average power dissipated in an a.c.circuit with an r.m.f current of 5A and voltage of 10v and phase angle of 60⁰
1
Expert's answer
2021-07-28T08:22:04-0400

The average power dissipated in an AC circuit 

"P_{\rm ave}=I_{\rm rms}V_{\rm rms}\cos\phi"

"P_{\rm ave}=5A*10V*\cos 60^{\circ}=25\:\rm W"